# NeRC: Neural Ranging Correction through Differentiable Moving Horizon Location Estimation

### v1

@@ 8. Related Work
We review related work about ranging error correction for GNSS, moving horizon estimation, and end-to-end structured learning.
Correcting Ranging Errors for GNSS: Correcting multipath and NLOS-induced pseudorange errors can significantly enhance GNSS positioning, but existing solutions often require 3D city models (Miura et al., 2015; Liu et al., 2018b) or additional sensors such as fisheye cameras or LiDAR (Wen et al., 2019; Bai et al., 2020), limiting widespread deployment. An alternative line of research uses deep learning to regress ranging errors directly from only raw GNSS measurements. While promising, state-of-the-art methods either depend on complex intermediate-loss estimation from fully labeled data (Sun et al., 2020; Zhang et al., 2021; Sun et al., 2023; Weng et al., 2024b) or learn via a basic WLS engine (Weng et al., 2024c; Hu et al., 2025), making them label-intensive or noise-sensitive. We address these limitations with a neural ranging correction framework that learns from widely available raw GNSS measurements through a differentiable, noise-resilient location estimator. Moreover, by incorporating basic 2D public maps, our method opens opportunities for leveraging ubiquitous unlabeled GNSS data.
@@
End-to-end Structured Learning: The learning paradigm that integrates data-driven neural modules and differentiable physical reasoning has been widely adopted in robotics (Li and Waslander, 2020; Wang et al., 2021, 2024; Cao et al., 2024), autonomous driving (Huang et al., 2023; Liu et al., 2025a), 3D vision (Chen et al., 2022; Valassakis and Garcia-Hernando, 2024), optical (Mildenhall et al., 2020) and radio frequency reconstruction (Zhao et al., 2023, 2024b), as well as other domains requiring physical or logical priors (Pineda et al., 2022). This paradigm has also been formalized under the high-level concept of Imperative Learning (Wang et al., 2025), with applications in path planning (Yang et al., 2023a), feature matching (Zhan et al., 2024), and simultaneous localization and mapping (SLAM) (Fu et al., 2024). Recent studies have brought this idea to GNSS localization, including end-to-end learning-based LOS/NLOS signal weighting (Xu et al., 2023a; Xu and Hsu, 2024) and ranging error correction (Weng et al., 2024c; Hu et al., 2025). In the context of ranging error regression, existing methods have so far relied solely on a baseline WLS localization engine and have been trained using fully labeled data with 3D ground truth (Weng et al., 2024c; Hu et al., 2025). In this paper, we address both limitations by introducing a differentiable, noise-resilient location estimator and enabling training with unlabeled data.
@@ 9. Conclusion, Limitations, and Future Work
This paper presents NeRC, a novel neural ranging correction framework designed to enhance GNSS localization for mobile devices. NeRC is trained through differentiable moving horizon location estimation, guided by a final location-based loss to ensure end-to-end optimization. Based on this learning pipeline, we propose a new paradigm that leverages unlabeled GNSS data by incorporating Euclidean distance supervision derived from publicly available maps. Extensive evaluations on both public and private benchmarks demonstrate the robustness and superiority of NeRC across rural and urban environments under both labeled and unlabeled settings. NeRC has reached the SOTA level. Furthermore, an edge-based field test reveals its promising potential for practical deployment.
The following limitations of NeRC merit further investigation:
(1) From centralized to distributed: NeRC is currently trained offline using pre-collected datasets on a high-performance workstation. To harness ubiquitous data and computational resources, training NeRC in a distributed way is a worthwhile direction (Du et al., 2023).
(2) From deep learning to pretraining: NeRC learns from unlabeled data under the assumption that we know the reference routes along which the data were collected. Exploiting more general crowdsourced unlabeled GNSS data with minimal prior knowledge could pave the way for pretraining GNSS foundation models (Smith et al., 2024).
(3) From fingerprinting to generalization: NeRC is currently designed to regress ranging errors of satellites observed in the same location during similar periods of days. Incorporating other sensing modalities shows potentials for generalizable error modeling (Weng et al., 2024a).
@@ Acknowledgements.
This paper was supported by Sponsor Nanyang Technological University https://www.ntu.edu.sg/ under the NTU Research Scholarship:˜Grant #R2015378.References
@@
- Axelrad et al. (2005) Penina Axelrad, Kristine Larson, and Brandon Jones. 2005. Use of the correct satellite repeat period to characterize and reduce site-specific multipath errors. In Proceedings of the 18th International Technical Meeting of the Satellite Division of the Institute of Navigation (ION GNSS 2005). Institute of Navigation, 2638–2648.
@@
- Bai et al. (2020) Xiwei Bai, Weisong Wen, and Li-ta Hsu. 2020. Using Sky-pointing fish-eye camera and LiDAR to aid GNSS single-point positioning in urban canyons. IET Intelligent Transport Systems 14, 8 (2020), 908–914.
@@
- Breu et al. (1995) Heinz Breu, Joseph Gil, David Kirkpatrick, and Michael Werman. 1995. Linear time Euclidean distance transform algorithms. IEEE Transactions on Pattern Analysis and Machine Intelligence 17, 5 (1995), 529–533.
@@
- Cao et al. (2024) Kun Cao, Xinhang Xu, Wanxin Jin, Karl H Johansson, and Lihua Xie. 2024. A differential dynamic programming framework for inverse reinforcement learning. arXiv preprint arXiv:2407.19902 (2024).
@@
- Chen et al. (2022) Hansheng Chen, Pichao Wang, Fan Wang, Wei Tian, Lu Xiong, and Hao Li. 2022. Epro-pnp: Generalized end-to-end probabilistic perspective-n-points for monocular object pose estimation. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 2781–2790.
@@
- Chen et al. (2024) Ying Chen, Sasamon Omoma, Hojung Kwon, Hazer Inaltekin, and Maria Gorlatova. 2024. Quantifying and Exploiting VR Frame Correlations: An Application of a Statistical Model for Viewport Pose. IEEE Transactions on Mobile Computing 23, 12 (2024), 11466–11482.
@@
- Ding and Chen (2025) Shiyi Ding and Ying Chen. 2025. RAG-VR: Leveraging Retrieval-Augmented Generation for 3D Question Answering in VR Environments. In 2025 IEEE Conference on Virtual Reality and 3D User Interfaces Abstracts and Workshops (VRW). IEEE, 131–136.
@@
- Du et al. (2023) Yao Du, Zehua Wang, Cyril Leung, and Victor CM Leung. 2023. Accelerating and securing blockchain-enabled distributed machine learning. IEEE Transactions on Mobile Computing 23, 6 (2023), 6712–6730.
@@
- Duan et al. (2024) Lin Duan, Ying Chen, Zhehan Qu, Megan McGrath, Erin Ehmke, and Maria Gorlatova. 2024. BiGuide: A bi-level data acquisition guidance for object detection on mobile devices. In 2024 23rd ACM/IEEE International Conference on Information Processing in Sensor Networks (IPSN). IEEE, 88–100.
@@
- Felzenszwalb and Huttenlocher (2012) Pedro F Felzenszwalb and Daniel P Huttenlocher. 2012. Distance transforms of sampled functions. Theory of computing 8, 1 (2012), 415–428.
@@
- Fiedler et al. (2020) Felix Fiedler, Dirk Baumbach, Anko Börner, and Sergio Lucia. 2020. A probabilistic moving horizon estimation framework applied to the visual-inertial sensor fusion problem. In 2020 European Control Conference (ECC). IEEE, 1009–1016.
@@
- Fu et al. (2020) Guoyu Michael Fu, Mohammed Khider, and Frank Van Diggelen. 2020. Android raw GNSS measurement datasets for precise positioning. In Proceedings of the 33rd international technical meeting of the satellite division of the Institute of Navigation (ION GNSS+ 2020). 1925–1937.
@@
- Fu et al. (2022) Michael Fu, Mohammed Khider, and Frank Van Diggelen. 2022. Summary and legacy of the smartphone decimeter challenge (sdc) 2022. In Proceedings of the 35th International Technical Meeting of the Satellite Division of The Institute of Navigation (ION GNSS+ 2022). 2301–2320.
@@
- Fu et al. (2024) Taimeng Fu, Shaoshu Su, Yiren Lu, and Chen Wang. 2024. islam: Imperative slam. IEEE Robotics and Automation Letters 9, 5 (2024), 4607–4614.
@@
- Han et al. (2019) Luxin Han, Fei Gao, Boyu Zhou, and Shaojie Shen. 2019. Fiesta: Fast incremental euclidean distance fields for online motion planning of aerial robots. In 2019 IEEE/RSJ International Conference on Intelligent Robots and Systems (IROS). IEEE, 4423–4430.
@@
- Haseltine and Rawlings (2005) Eric L Haseltine and James B Rawlings. 2005. Critical evaluation of extended Kalman filtering and moving-horizon estimation. Industrial & engineering chemistry research 44, 8 (2005), 2451–2460.
@@
- He et al. (2015) Kaiming He, Xiangyu Zhang, Shaoqing Ren, and Jian Sun. 2015. Delving deep into rectifiers: Surpassing human-level performance on imagenet classification. In Proceedings of the IEEE international conference on computer vision. 1026–1034.
@@
- Hsu (2018) Li-Ta Hsu. 2018. Analysis and modeling GPS NLOS effect in highly urbanized area. GPS Solutions 22, 1 (2018), 7. doi:10.1007/s10291-017-0667-9
- Hsu et al. (2023) Li-Ta Hsu, Feng Huang, Hoi-Fung Ng, Guohao Zhang, Yihan Zhong, Xiwei Bai, and Weisong Wen. 2023. Hong Kong UrbanNav: An open-source multisensory dataset for benchmarking urban navigation algorithms. NAVIGATION: Journal of the Institute of Navigation 70, 4 (2023).
- Hu et al. (2025) Runzhi Hu, Penghui Xu, Yihan Zhong, and Weisong Wen. 2025. pyrtklib: An open-source package for tightly coupled deep learning and GNSS integration for positioning in urban canyons. IEEE Transactions on Intelligent Transportation Systems (2025).
- Hu et al. (2024) Tianyi Hu, Tim Scargill, Fan Yang, Ying Chen, Guohao Lan, and Maria Gorlatova. 2024. SEESys: Online pose error estimation system for visual SLAM. In Proceedings of the 22nd ACM Conference on Embedded Networked Sensor Systems. 322–335.
@@
- Jazwinski (1970) AH Jazwinski. 1970. Stochastic processes and filtering theory. ACADEMIC PRESS, INC.,.
- Kang et al. (2024) Jiarong Kang, Yi Wang, and Xiaobin Xiong. 2024. Fast decentralized state estimation for legged robot locomotion via EKF and MHE. IEEE Robotics and Automation Letters (2024).
- Kanhere et al. (2022) Ashwin V Kanhere, Shubh Gupta, Akshay Shetty, and Grace Gao. 2022. Improving gnss positioning using neural-network-based corrections. NAVIGATION: Journal of the Institute of Navigation 69, 4 (2022).
- Kaplan and Hegarty (2017) Elliott D Kaplan and Christopher Hegarty. 2017. Understanding GPS/GNSS: Principles and Applications. Artech house.
- Li and Waslander (2020) Chunshang Li and Steven L Waslander. 2020. Towards end-to-end learning of visual inertial odometry with an EKF. In 2020 17th Conference on Computer and Robot Vision (CRV). IEEE Computer Society, 190–197.
@@
- Liu et al. (2025a) Haochen Liu, Zhiyu Huang, Wenhui Huang, Haohan Yang, Xiaoyu Mo, and Chen Lv. 2025a. Hybrid-prediction integrated planning for autonomous driving. IEEE Transactions on Pattern Analysis and Machine Intelligence (2025).
@@
- Liu et al. (2025b) Peng Liu, Honglei Qin, Jun Lu, Muyuan Jiang, Yong Liang Guan, and Chau Yuen. 2025b. Attitude Bound-Constrained Factor Graph Optimization for GNSS and IMU Fusion Positioning in Bumpy Scenarios. IEEE Transactions on Instrumentation and Measurement (2025).
@@
- Mohanty and Gao (2023) Adyasha Mohanty and Grace Gao. 2023. Learning GNSS positioning corrections for smartphones using graph convolution neural networks. NAVIGATION: Journal of the Institute of Navigation 70, 4 (2023).
- Morton et al. (2021) Y Jade Morton, Frank van Diggelen, James J Spilker Jr, Bradford W Parkinson, Sherman Lo, and Grace Gao. 2021. Position, navigation, and timing technologies in the 21st century: Integrated satellite navigation, sensor systems, and civil applications, volume 1. John Wiley & Sons.
@@
- Rao and Rawlings (2000) Christopher V Rao and James B Rawlings. 2000. Nonlinear moving horizon state estimation. In Nonlinear Model Predictive Control. Springer, 45–69.
@@
- Ratliff et al. (2009) Nathan Ratliff, Matt Zucker, J Andrew Bagnell, and Siddhartha Srinivasa. 2009. CHOMP: Gradient optimization techniques for efficient motion planning. In 2009 IEEE International Conference on Robotics and Automation. IEEE, 489–494.
- Sarlin et al. (2023) Paul-Edouard Sarlin, Daniel DeTone, Tsun-Yi Yang, Armen Avetisyan, Julian Straub, Tomasz Malisiewicz, Samuel Rota Bulo, Richard Newcombe, Peter Kontschieder, and Vasileios Balntas. 2023. Orienternet: Visual localization in 2d public maps with neural matching. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 21632–21642.
- Scargill et al. (2023) Tim Scargill, Ying Chen, Tianyi Hu, and Maria Gorlatova. 2023. SiTAR: Situated trajectory analysis for in-the-wild pose error estimation. In 2023 IEEE International Symposium on Mixed and Augmented Reality (ISMAR). IEEE, 283–292.
@@
- Spilker Jr et al. (1996) James J Spilker Jr, Penina Axelrad, Bradford W Parkinson, and Per Enge. 1996. Global positioning system: theory and applications, volume I. American Institute of Aeronautics and Astronautics.
@@
- Sun et al. (2025) Rui Sun, Qi Sheng, Qi Cheng, Xiaotong Shang, and Washington Yotto Ochieng. 2025. 3D Grid-Based Resilient Pseudorange Error Prediction for Adaptive GNSS/IMU Integrated Navigation in Urban Areas. IEEE Internet of Things Journal (2025).
@@
- Tenny and Rawlings (2002) Matthew J Tenny and James B Rawlings. 2002. Efficient moving horizon estimation and nonlinear model predictive control. In Proceedings of the 2002 American Control Conference (IEEE Cat. No. CH37301), Vol. 6. IEEE, 4475–4480.
- Valassakis and Garcia-Hernando (2024) Eugene Valassakis and Guillermo Garcia-Hernando. 2024. Handdgp: Camera-space hand mesh prediction with differentiable global positioning. In European Conference on Computer Vision. Springer, 479–496.
@@
- Wang et al. (2021) Bingheng Wang, Zhengtian Ma, Shupeng Lai, Lin Zhao, and Tong Heng Lee. 2021. Differentiable moving horizon estimation for robust flight control. In 2021 60th IEEE Conference on Decision and Control (CDC). IEEE, 3563–3568.

### v2

@@ 8. Related Work
Correcting Ranging Errors for GNSS: Correcting multipath and NLOS-induced pseudorange errors can enhance GNSS positioning, but existing solutions often require 3D city models (Miura et al., 2015; Liu et al., 2018b) or additional sensors such as fisheye cameras or LiDAR (Wen et al., 2019; Bai et al., 2020), limiting widespread deployment. An alternative line of research uses deep learning to regress ranging errors directly from only raw GNSS measurements. While promising, the related methods either depend on complex intermediate-loss estimation from fully labeled data (Sun et al., 2020; Zhang et al., 2021; Sun et al., 2023; Weng et al., 2024b), making them label-intensive and less automatic. We address these limitations by designing a neural ranging correction framework that is trained and operates end-to-end.
@@
End-to-end Structured Learning: The learning paradigm that integrates data-driven neural modules and differentiable physical reasoning has been widely adopted in robotics (Li and Waslander, 2020; Wang et al., 2021, 2024; Cao et al., 2025), autonomous driving (Huang et al., 2023; Liu et al., 2025a), 3D vision (Chen et al., 2025; Valassakis and Garcia-Hernando, 2025), optical (Mildenhall et al., 2020) and radio frequency reconstruction (Zhao et al., 2023, 2024b), as well as other domains requiring physical or logical priors (Pineda et al., 2022). This paradigm has also been formalized under the high-level concept of Imperative Learning (Wang et al., 2025; Yang et al., 2023; Zhan et al., 2024; Fu et al., 2024). Recent studies have brought this idea to GNSS localization, including end-to-end learning-based LOS/NLOS signal weighting (Xu et al., 2023a; Xu and Hsu, 2024) and ranging error correction (Weng et al., 2024c; Hu et al., 2025). Regarding ranging error regression, existing methods have so far relied solely on a baseline WLS localization engine and have been trained using fully labeled data (Weng et al., 2024c; Hu et al., 2025). This paper addresses both limitations by introducing a differentiable, noise-resilient location estimator and incorporating basic 2D public maps to enable training with ubiquitous unlabeled GNSS data.
@@ 9. Conclusion, Limitations, and Future Work
This paper presents NeRC, a novel neural ranging correction framework that enhances GNSS localization for mobile devices. NeRC is trained through differentiable MHE, guided by location-based losses to ensure end-to-end optimization. Based on this, we propose a new training paradigm that can leverage unlabeled GNSS data by incorporating Euclidean-distance supervision derived from publicly available maps. Extensive evaluations across both public and private benchmarks demonstrate the robustness and superiority of NeRC across rural and urban environments under both labeled and unlabeled settings. Furthermore, an edge-based field test reveals its promising potential for practical deployment. The following limitations of NeRC merit further investigation:
(1) From centralized to distributed: Enabling onboard training and inference with data from heterogeneous devices is important for NeRC’s distributed deployment (Du et al., 2023; Xu et al., 2023b; Du et al., 2025). It is also necessary to systematically investigate the spatial-temporal granularity at which NeRC models should be trained and distributed for large-scale use.
(2) From deep learning to pretraining: NeRC learns from unlabeled data under the assumption that we know data collection routes. Exploiting crowdsourced data with minimal prior knowledge could pave the way for GNSS foundation models (Smith et al., 2024).
(3) From fingerprinting to generalization: NeRC is currently designed to regress ranging errors of satellites observed in the same region during similar periods of days. Incorporating other sensing modalities shows potential for generalizable error modeling (Weng et al., 2024a).
@@ Acknowledgements.
This work was supported by Sponsor Nanyang Technological University https://www.ntu.edu.sg/ under the NTU Research Scholarship:˜Grant #R2015378. Kun Cao was supported by the National Natural Science Foundation of China under Grant 62088101, 62503367, Shanghai Municipal Science and Technology Major Project (No. 2021SHZDZX0100).References
@@
- Axelrad et al. (2005) Penina Axelrad, Kristine Larson, and Brandon Jones. 2005. Use of the Correct Satellite Repeat Period to Characterize and Reduce Site-Specific Multipath Errors. In Proceedings of the 18th International Technical Meeting of the Satellite Division of the Institute of Navigation (ION GNSS 2005). Institute of Navigation, Long Beach, CA, 2638–2648.
@@
- Bai et al. (2020) Xiwei Bai, Weisong Wen, and Li-ta Hsu. 2020. Using Sky-Pointing Fish-Eye Camera and LiDAR to Aid GNSS Single-Point Positioning in Urban Canyons. IET Intelligent Transport Systems 14, 8 (2020), 908–914. doi:10.1049/iet-its.2019.0587
@@
- Breu et al. (1995) Heinz Breu, Joseph Gil, David Kirkpatrick, and Michael Werman. 1995. Linear Time Euclidean Distance Transform Algorithms. IEEE Transactions on Pattern Analysis and Machine Intelligence 17, 5 (1995), 529–533. doi:10.1109/34.391389
@@
- Cao et al. (2025) Kun Cao, Xinhang Xu, Wanxin Jin, Karl H Johansson, and Lihua Xie. 2025. A Differential Dynamic Programming Framework for Inverse Reinforcement Learning. IEEE Transactions on Robotics (2025). doi:10.1109/TRO.2025.3623769
@@
- Chen et al. (2025) Hansheng Chen, Wei Tian, Pichao Wang, Fan Wang, Lu Xiong, and Hao Li. 2025. EPro-PnP: Generalized End-to-End Probabilistic Perspective-n-Points for Monocular Object Pose Estimation. IEEE Transactions on Pattern Analysis and Machine Intelligence 47, 11 (2025), 9413–9425. doi:10.1109/TPAMI.2024.3354997
@@
- Chen et al. (2024) Ying Chen, Sasamon Omoma, Hojung Kwon, Hazer Inaltekin, and Maria Gorlatova. 2024. Quantifying and Exploiting VR Frame Correlations: An Application of a Statistical Model for Viewport Pose. IEEE Transactions on Mobile Computing 23, 12 (2024), 11466–11482. doi:10.1109/TMC.2024.3399770
@@
- Ding and Chen (2025) Shiyi Ding and Ying Chen. 2025. RAG-VR: Leveraging Retrieval-Augmented Generation for 3D Question Answering in VR Environments. In 2025 IEEE Conference on Virtual Reality and 3D User Interfaces Abstracts and Workshops (VRW). IEEE, Saint Malo, France, 131–136. doi:10.1109/VRW66409.2025.00034
- Du et al. (2025) Yao Du, Cyril Leung, Zehua Wang, Xiaoxiao Li, and Victor Leung. 2025. Decentralized Model Selection for Test-Time Adaptation in Heterogeneous Connected Systems. ACM Transactions on the Web (2025). doi:10.1145/3764936
@@
- Du et al. (2023) Yao Du, Zehua Wang, Cyril Leung, and Victor CM Leung. 2023. Accelerating and Securing Blockchain-Enabled Distributed Machine Learning. IEEE Transactions on Mobile Computing 23, 6 (2023), 6712–6730. doi:10.1109/TMC.2023.3325334
@@
- Duan et al. (2024) Lin Duan, Ying Chen, Zhehan Qu, Megan McGrath, Erin Ehmke, and Maria Gorlatova. 2024. BiGuide: A Bi-level Data Acquisition Guidance for Object Detection on Mobile Devices. In 2024 23rd ACM/IEEE International Conference on Information Processing in Sensor Networks (IPSN). IEEE, Hong Kong, China, 88–100. doi:10.1109/IPSN61024.2024.00012
@@
- Felzenszwalb and Huttenlocher (2012) Pedro F Felzenszwalb and Daniel P Huttenlocher. 2012. Distance Transforms of Sampled Functions. Theory of computing 8, 1 (2012), 415–428. doi:10.4086/toc.2012.v008a019
@@
- Fiedler et al. (2020) Felix Fiedler, Dirk Baumbach, Anko Börner, and Sergio Lucia. 2020. A Probabilistic Moving Horizon Estimation Framework Applied to the Visual-Inertial Sensor Fusion Problem. In 2020 European Control Conference (ECC). IEEE, St. Petersburg, Russia, 1009–1016. doi:10.23919/ECC51009.2020.9143645
@@
- Fu et al. (2020) Guoyu Michael Fu, Mohammed Khider, and Frank Van Diggelen. 2020. Android Raw GNSS Measurement Datasets for Precise Positioning. In Proceedings of the 33rd International Technical Meeting of the Satellite Division of the Institute of Navigation (ION GNSS+ 2020). Institute of Navigation, Virtual Program, 1925–1937. doi:10.33012/2020.17628
@@
- Fu et al. (2022) Michael Fu, Mohammed Khider, and Frank Van Diggelen. 2022. Summary and Legacy of the Smartphone Decimeter Challenge (SDC) 2022. In Proceedings of the 35th International Technical Meeting of the Satellite Division of The Institute of Navigation (ION GNSS+ 2022). Institute of Navigation, Denver, Colorado, 2301–2320. doi:10.33012/2022.18379
@@
- Fu et al. (2024) Taimeng Fu, Shaoshu Su, Yiren Lu, and Chen Wang. 2024. iSLAM: Imperative SLAM. IEEE Robotics and Automation Letters 9, 5 (2024), 4607–4614. doi:10.1109/LRA.2024.3382533
@@
- Han et al. (2019) Luxin Han, Fei Gao, Boyu Zhou, and Shaojie Shen. 2019. FIESTA: Fast Incremental Euclidean Distance Fields for Online Motion Planning of Aerial Robots. In 2019 IEEE/RSJ International Conference on Intelligent Robots and Systems (IROS). IEEE, Macau, China, 4423–4430. doi:10.1109/IROS40897.2019.8968199
@@
- Haseltine and Rawlings (2005) Eric L Haseltine and James B Rawlings. 2005. Critical Evaluation of Extended Kalman Filtering and Moving-Horizon Estimation. Industrial & engineering chemistry research 44, 8 (2005), 2451–2460. doi:10.1021/ie034308l
@@
- He et al. (2015) Kaiming He, Xiangyu Zhang, Shaoqing Ren, and Jian Sun. 2015. Delving Deep into Rectifiers: Surpassing Human-Level Performance on Imagenet Classification. In Proceedings of the IEEE International Conference on Computer Vision. IEEE, Los Alamitos, CA, USA, 1026–1034. doi:10.1109/ICCV.2015.123
@@
- Hsu (2018) Li-Ta Hsu. 2018. Analysis and Modeling GPS NLOS Effect in Highly Urbanized Area. GPS Solutions 22, 1 (2018), 1–12. doi:10.1007/s10291-017-0667-9
- Hsu et al. (2023) Li-Ta Hsu, Feng Huang, Hoi-Fung Ng, Guohao Zhang, Yihan Zhong, Xiwei Bai, and Weisong Wen. 2023. Hong Kong UrbanNav: An open-source multisensory dataset for benchmarking urban navigation algorithms. NAVIGATION: Journal of the Institute of Navigation 70, 4 (2023), 1–29. doi:10.33012/navi.602
- Hu et al. (2025) Runzhi Hu, Penghui Xu, Yihan Zhong, and Weisong Wen. 2025. pyrtklib: An Open-Source Package for Tightly Coupled Deep Learning and GNSS Integration for Positioning in Urban Canyons. IEEE Transactions on Intelligent Transportation Systems 26, 7 (2025), 10652–10662. doi:10.1109/TITS.2025.3552691
- Hu et al. (2024) Tianyi Hu, Tim Scargill, Fan Yang, Ying Chen, Guohao Lan, and Maria Gorlatova. 2024. SEESys: Online pose error estimation system for visual SLAM. In Proceedings of the 22nd ACM Conference on Embedded Networked Sensor Systems (Hangzhou, China) (SenSys ’24). Association for Computing Machinery, New York, NY, USA, 322–335. doi:10.1145/3666025.3699341
@@
- Jazwinski (1970) Andrew H. Jazwinski. 1970. Stochastic Processes and Filtering Theory. Analytical Mechanics Associates, Inc. Seabrook, Maryland.
- Kang et al. (2024) Jiarong Kang, Yi Wang, and Xiaobin Xiong. 2024. Fast Decentralized State Estimation for Legged Robot Locomotion via EKF and MHE. IEEE Robotics and Automation Letters 9, 12 (2024), 10914–10921. doi:10.1109/LRA.2024.3483043
- Kanhere et al. (2022) Ashwin V Kanhere, Shubh Gupta, Akshay Shetty, and Grace Gao. 2022. Improving GNSS Positioning Using Neural-Network-Based Corrections. NAVIGATION: Journal of the Institute of Navigation 69, 4 (2022), 1–20. doi:10.33012/navi.548
- Kaplan and Hegarty (2017) Elliott D Kaplan and Christopher Hegarty. 2017. Understanding GPS/GNSS: Principles and Applications. Artech house, Boston — London.
- Li and Waslander (2020) Chunshang Li and Steven L Waslander. 2020. Towards End-to-End Learning of Visual Inertial Odometry with an EKF. In 2020 17th Conference on Computer and Robot Vision (CRV). IEEE, Ottawa, ON, Canada, 190–197. doi:10.1109/CRV50864.2020.00033
@@
- Liu et al. (2025a) Haochen Liu, Zhiyu Huang, Wenhui Huang, Haohan Yang, Xiaoyu Mo, and Chen Lv. 2025a. Hybrid-Prediction Integrated Planning for Autonomous Driving. IEEE Transactions on Pattern Analysis and Machine Intelligence 47, 4 (2025), 2597–2614. doi:10.1109/TPAMI.2025.3526936
@@
- Liu et al. (2025b) Peng Liu, Honglei Qin, Jun Lu, Muyuan Jiang, Yong Liang Guan, and Chau Yuen. 2025b. Attitude Bound-Constrained Factor Graph Optimization for GNSS and IMU Fusion Positioning in Bumpy Scenarios. IEEE Transactions on Instrumentation and Measurement 74 (2025), 1–11. doi:10.1109/TIM.2025.3586270
@@
- Mohanty and Gao (2023) Adyasha Mohanty and Grace Gao. 2023. Learning GNSS Positioning Corrections for Smartphones Using Graph Convolution Neural Networks. NAVIGATION: Journal of the Institute of Navigation 70, 4 (2023), 1–15. doi:10.33012/navi.622
- Morton et al. (2021) Y Jade Morton, Frank van Diggelen, James J Spilker Jr, Bradford W Parkinson, Sherman Lo, and Grace Gao. 2021. Position, Navigation, and Timing Technologies in the 21st Century: Integrated Satellite Navigation, Sensor Systems, and Civil Applications. John Wiley & Sons, Inc., Hoboken, New Jersey. doi:10.1002/9781119458449
@@
- Rao and Rawlings (2000) Christopher V Rao and James B Rawlings. 2000. Nonlinear Moving Horizon State Estimation. In Nonlinear Model Predictive Control. Birkhäuser, Basel, 45–69. doi:10.1007/978-3-0348-8407-5_3
@@
- Ratliff et al. (2009) Nathan Ratliff, Matt Zucker, J Andrew Bagnell, and Siddhartha Srinivasa. 2009. CHOMP: Gradient Optimization Techniques for Efficient Motion Planning. In 2009 IEEE International Conference on Robotics and Automation. IEEE, Kobe, Japan, 489–494. doi:10.1109/ROBOT.2009.5152817
- Sarlin et al. (2023) Paul-Edouard Sarlin, Daniel DeTone, Tsun-Yi Yang, Armen Avetisyan, Julian Straub, Tomasz Malisiewicz, Samuel Rota Bulo, Richard Newcombe, Peter Kontschieder, and Vasileios Balntas. 2023. Orienternet: Visual Localization in 2D Public Maps with Neural Matching. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. IEEE, Vancouver, BC, Canada, 21632–21642. doi:10.1109/CVPR52729.2023.02072
- Scargill et al. (2023) Tim Scargill, Ying Chen, Tianyi Hu, and Maria Gorlatova. 2023. SiTAR: Situated Trajectory Analysis for In-the-Wild Pose Error Estimation. In 2023 IEEE International Symposium on Mixed and Augmented Reality (ISMAR). IEEE, Sydney, Australia, 283–292. doi:10.1109/ISMAR59233.2023.00043
@@
- Spilker Jr et al. (1996) James J Spilker Jr, Penina Axelrad, Bradford W Parkinson, and Per Enge. 1996. Global Positioning System: Theory and Applications, Volume I. American Institute of Aeronautics and Astronautics, Inc., 370 L’Enfant Promenade, SW, Washington, DC 20024-2518. doi:10.2514/4.866388
@@
- Sun et al. (2025) Rui Sun, Qi Sheng, Qi Cheng, Xiaotong Shang, and Washington Yotto Ochieng. 2025. 3D Grid-Based Resilient Pseudorange Error Prediction for Adaptive GNSS/IMU Integrated Navigation in Urban Areas. IEEE Internet of Things Journal 12, 12 (2025), 19264–19279. doi:10.1109/JIOT.2025.3533077
@@
- Tenny and Rawlings (2002) Matthew J Tenny and James B Rawlings. 2002. Efficient Moving Horizon Estimation and Nonlinear Model Predictive Control. In Proceedings of the 2002 American Control Conference, Vol. 6. IEEE, Anchorage, AK, USA, 4475–4480. doi:10.1109/ACC.2002.1025355
- Valassakis and Garcia-Hernando (2025) Eugene Valassakis and Guillermo Garcia-Hernando. 2025. Handdgp: Camera-Space Hand Mesh Prediction with Differentiable Global Positioning. In Computer Vision – ECCV 2024. Springer Nature Switzerland, Cham, 479–496. doi:10.1007/978-3-031-72920-1_27
@@
- Wang et al. (2021) Bingheng Wang, Zhengtian Ma, Shupeng Lai, Lin Zhao, and Tong Heng Lee. 2021. Differentiable Moving Horizon Estimation for Robust Flight Control. In 2021 60th IEEE Conference on Decision and Control (CDC). IEEE, Austin, TX, USA, 3563–3568. doi:10.1109/CDC45484.2021.9683173